# Decoherence of quantum superpositions by Reissner–Nordström black holes

### v1

@@ Abstract
Recently, it was shown by Danielson-Satishchandran-Wald (DSW) that for the massive or charged body in a quantum spatial separated superposition state, the presence of a black hole can decohere the superposition inevitably towards capturing the radiation of soft photons or gravitons Danielson:2022tdw ; Danielson:2022sga ; Danielson:2024yru . In this work, we study the DSW decoherence effect for the static charged body in the Reissner–Nordström black holes. By calculating the decohering rate for this case, it is shown that the superposition is decohered by the low frequency photons that propagate through the black hole horizon. For the extremal Reissner–Nordström black hole, the decoherence of quantum superposition is completely suppressed due to the black hole Meissner effect.
@@ 1 Introduction
In this work, we study the DSW decoherence effect for a static charged body in the rn black holes. Our calculation is limited to the case of electromagnetic radiation from the charged particle in a spatial separated quantum superposition state. By calculating the decohering rate for the current case, it is shown that the superposition is decohered by the low frequency photons that propagate through the black hole horizon. In other words, we provide a detailed calculation demonstrating that the quantum superposition is violated due to the entanglement between the black hole and the charged particle, which arises from the black hole absorbing the entangling photon emitted by the charged particle. We also discuss the case of the extremal Reissner–Nordström black hole. It is shown that the decoherence of quantum superposition is completely suppressed due to the black hole Meissner effect Wald:1974np ; King:1975tt ; Bicak:1980du ; Bini:2008iwy .
This paper is arranged as follows. In Sec.2, we briefly review the setup of the DSW Gedankenexperiment for decoherence of quantum superposition. In Sec.3, we discuss how the decoherence rises from the entangling photon radiation. In Sec.4, we make an estimation of the entangling photon number in the Reissner–Nordström black hole spacetime. In Sec.5, we calculate the decohering flux and give an exact expression for the decoherence rate. The conclusion and the discussion are presented in the last section.
@@ 3 Decoherence from the entanglement photons
After the time at which the particle is recombined, the ”out” radiation field can also be defined analogous to Eq.(2)
@@
By a simple estimation, it can be shown that the decoherence from the future null infinity is zero. This is caused by the fact that the decay of the gauge potential asymptotically like $\frac{1}{r}$ in the spherically symmetric spacetime Danielson:2021egj . Then the contribution from the future null infinity to the entangling photon number is asymptotical $\frac{1}{r^{2}}$, which is zero when taking the limit $r\rightarrow+\infty$. Therefore, to evaluate the entangling photon number, one can safely calculate the inner product on the future horizon. In the following, our aim is to evaluate the expected number of entangling photons at late times for the static charged particle in the Reissner–Nordström black hole.
4 Estimation of entangling photon number in Reissner–Nordström black holes
The Reissner–Nordström black hole is described by the metric
@@ 4 Estimation of entangling photon number in Reissner–Nordström black holes
Using the Kruskal-like coordinates (19), the Reissner–Nordström metric has the form of
@@
which means the entangling photon number is linearly growing with the Killing time. This results indicates that the quantum superposition state hold by the experimenter will be decohered after sufficiently long time by radiation the low frequency soft photon into the black hole horizon. However, in the extremal black hole limit, where $\kappa_{+}\rightarrow 0$, the entangling photon number will be sufficiently suppressed and the coherence of the spatial separated quantum superposition will be maintained. Thus far, our final task is to compute the exact expression of the decohering flux given in Eq.(27) for the Reissner–Nordström black holes.
@@ 5 Calculation of decohering flux
To compute the exact expression of the decohering flux, we assume that the left and the right branches of the charge currents persist forever. Then the left and the right branches of the charge current corresponding to the spatial quantum superposition state can be treated as the static sources. Therefore, we firstly study the static electromagnetic field generated by the classical static charged source in Reissner–Nordström black hole.
@@
where $q$ is the electric charge and $u^{\mu}=\frac{dx^{\mu}}{d\tau}$ is the four velocity of the charged particle. For a static charge on $z$ axis of a Reissner–Nordström black hole, we have
@@
It is easy to see that the solution is also valid for the extremal Reissner–Nordström black hole with $M=Q$. In this case the Maxwell equation (30) is formally identical to Laplace’s equation in Minkowski space.
@@
where $x^{A}=\{\theta,\phi\}$ is the transverse coordinates, $q_{AB}=\textrm{diag}\{r_{+}^{2},r_{+}^{2}\sin^{2}\theta\}$ is the transverse metric on the horizon, $\epsilon_{AB}$ is the two-dimensional area element, and ${}^{*}F_{AB}$ is the dual electromagnetic tensor.
For the Reissner–Nordström black hole spacetime, the Maxwell equations on the horizon $r=r_{+}$ are explicitly given by
$\frac{1}{\sqrt{q}}\partial_{\theta}\left(\sqrt{q}q^{\theta\theta}A^{c}_{\theta}\right)$ $=-\frac{1}{2}\epsilon^{AB}\ \!{}^{*}\!F^{c}_{AB}\;,$ (35)
$\frac{1}{\sqrt{q}}\partial_{\theta}A^{c}_{\bar{\phi}}$ $=\frac{1}{2}\epsilon^{AB}F^{c}_{AB}\;.$ (36)
By assuming that $A_{A}^{c}$ is independent of the angular coordinate $\phi$, the integrals can be done in closed form. After some calculation, one can find the nonvanishing component of the ”Coulomb” gauge potential on the horizon as
$A_{\theta}^{c}=\frac{-qr_{+}(r_{+}-M)\sin\theta}{r_{0}\left[r_{0}-M-(r_{+}-M)\cos\theta\right]}\;,$ (37)
where an integral constant $-\frac{qr_{+}}{r_{0}}\frac{r_{0}-M}{r_{+}-M}$ has been selected to eliminate the divergence on the pole. The above equation (37) refers to the Coulomb field generated by a single static point charge $q$ in the Reissner–Nordström black hole spacetime. It is clear that the gauge potential vanishes at the extremal limit $r_{+}\rightarrow M$, which is an indication of the black hole Meissner effect Wald:1974np ; King:1975tt ; Bicak:1980du ; Bini:2008iwy .
@@
$\Delta A_{\theta}=\frac{pr_{+}((r_{+}-M))\left(M-2r_{0}+(r_{+}-M)\cos\theta\right)\sin\theta}{r_{0}^{2}\sqrt{g_{rr}}|_{r=r_{0}}\left[r_{0}-M-(r_{+}-M)\cos\theta\right]^{2}}\;.$ (39)
@@
This result along with Eq.(28) gives the exact expression for the expected number of the entangling photon generated by the charged particle in the spatial separated quantum superposition state in the Reissner–Nordström black hole.
@@ 6 Conclusion and discussion
In summary, we have studied the DSW decoherence effect for a static charged particle in the rn black holes. The explicit expression for the decoherence is obtained by evaluating the expected number of the entangling photon at late times. Our calculation is limited to the case of electromagnetic radiation from the charged particle in a spatial separated quantum superposition state. It is shown that the quantum superposition is decohered by the low frequency photons that propagate through the black hole horizon. We also discuss the case of the extremal Reissner–Nordström black hole. It is shown that the decoherence of quantum superposition is completely suppressed due to the black hole Meissner effect.

### v2

# Decoherence of quantum superpositions by Reissner-Nordström black holes

@@ Abstract
Recently, it was shown by Danielson-Satishchandran-Wald (DSW) that for the massive or charged body in a quantum spatial separated superposition state, the presence of a black hole can decohere the superposition inevitably towards capturing the radiation of soft photons or gravitons Danielson:2022tdw ; Danielson:2022sga ; Danielson:2024yru . In this work, we study the DSW decoherence effect for the static charged body in the Reissner-Nordström black holes. By calculating the decohering rate for this case, it is shown that the superposition is decohered by the low frequency photons that propagate through the black hole horizon. For the extremal Reissner-Nordström black hole, the decoherence of quantum superposition is completely suppressed due to the black hole Meissner effect. It is also found that the decoherence effect caused by the Reissner-Nordström black hole is equivalent to that of an ordinary matter system with the same size and charge.
@@ 1 Introduction
In this work, we study the DSW decoherence effect for a static charged body in the rn black holes. Our calculation is limited to the case of electromagnetic radiation from the charged particle in a spatial separated quantum superposition state. By calculating the decohering rate for the current case, it is shown that the superposition is decohered by the low frequency photons that propagate through the black hole horizon. In other words, we provide a detailed calculation demonstrating that the quantum superposition is violated due to the entanglement between the black hole and the charged particle, which arises from the black hole absorbing the entangling photon emitted by the charged particle. We also discuss the case of the extremal Reissner-Nordström black hole. It is shown that the decoherence of quantum superposition is completely suppressed due to the black hole Meissner effect Wald:1974np ; King:1975tt ; Bicak:1980du ; Bini:2008iwy . At last, we compare the decoherence effect due to the Reissner-Nordström black hole with that of the ordinary material system and find that the decoherence effect caused by the Reissner-Nordström black hole is equivalent to that of an ordinary matter system with the same size and charge.
This paper is arranged as follows. In Sec.2, we briefly review the setup of the DSW gedankenexperiment for decoherence of quantum superposition. In Sec.3, we discuss how the decoherence rises from the entangling photon radiation. In Sec.4, we make an estimation of the entangling photon number in the Reissner-Nordström black hole spacetime. In Sec.5, we calculate the decohering flux and give an exact expression for the decoherence rate. In Sec.6, we compare the decoherence effect due to the Reissner-Nordström black hole with that of the ordinary material system. The conclusion and the discussion are presented in the last section.
@@ 3 Decoherence from the entanglement photons
After the time at which the particle is recombined, the “out” radiation field can also be defined analogous to Eq.(2)
@@
By a simple estimation, it can be shown that the decoherence from the future null infinity is zero. This is caused by the fact that the decay of the gauge potential asymptotically like $\frac{1}{r}$ in the spherically symmetric spacetime Danielson:2021egj . Then the contribution from the future null infinity to the entangling photon number is asymptotical $\frac{1}{r^{2}}$, which is zero when taking the limit $r\rightarrow+\infty$. Therefore, to evaluate the entangling photon number, one can safely calculate the inner product on the future horizon. In the following, our aim is to evaluate the expected number of entangling photons at late times for the static charged particle in the Reissner-Nordström black hole.
4 Estimation of entangling photon number in Reissner-Nordström black holes
The Reissner-Nordström black hole is described by the metric
@@ 4 Estimation of entangling photon number in Reissner-Nordström black holes
Using the Kruskal-like coordinates (19), the Reissner-Nordström metric has the form of
@@
which means the entangling photon number is linearly growing with the Killing time. This results indicates that the quantum superposition state hold by the experimenter will be decohered after sufficiently long time by radiation the low frequency soft photon into the black hole horizon. However, in the extremal black hole limit, where $\kappa_{+}\rightarrow 0$, the entangling photon number will be sufficiently suppressed and the coherence of the spatial separated quantum superposition will be maintained. Thus far, our final task is to compute the exact expression of the decohering flux given in Eq.(27) for the Reissner-Nordström black holes.
@@ 5 Calculation of decohering flux
To compute the exact expression of the decohering flux, we assume that the left and the right branches of the charge currents persist forever. Then the left and the right branches of the charge current corresponding to the spatial quantum superposition state can be treated as the static sources. Therefore, we firstly study the static electromagnetic field generated by the classical static charged source in Reissner-Nordström black hole.
@@
where $q$ is the electric charge and $u^{\mu}=\frac{dx^{\mu}}{d\tau}$ is the four velocity of the charged particle. For a static charge on $z$ axis of a Reissner-Nordström black hole, we have
@@
It is easy to see that the solution is also valid for the extremal Reissner-Nordström black hole with $M=Q$. In this case the Maxwell equation (30) is formally identical to Laplace’s equation in Minkowski space.
@@
where $x^{A}=\{\theta,\phi\}$ is the transverse coordinates, $q_{AB}=\textrm{diag}\{r_{+}^{2},r_{+}^{2}\sin^{2}\theta\}$ is the transverse metric on the horizon, $\epsilon_{AB}$ is the two-dimensional area element, and ${}^{*}F_{AB}$ is the dual electromagnetic tensor. It should be noted that the right hand sides of Eq.(33) and Eq.(34) are just the radial electric and magnetic fields, which are clearly gauge invariant.
For the Reissner-Nordström black hole spacetime, the Maxwell equations on the horizon $r=r_{+}$ are explicitly given by
$\frac{1}{\sqrt{q}}\partial_{\theta}\left(\sqrt{q}q^{\theta\theta}A^{c}_{\theta}\right)$ $=\frac{1}{2}\epsilon^{AB}\ \!{}^{*}\!F^{c}_{AB}\;,$ (35)
$\frac{1}{\sqrt{q}}\partial_{\theta}A^{c}_{\phi}$ $=\frac{1}{2}\epsilon^{AB}F^{c}_{AB}\;.$ (36)
By assuming that $A_{A}^{c}$ is independent of the angular coordinate $\phi$, the integrals can be done in closed form. After some calculation (see Appendix A), one can find the nonvanishing component of the “Coulomb” gauge potential on the horizon as
$A_{\theta}^{c}=\frac{qr_{+}(r_{+}-M)\sin\theta}{r_{0}\left[r_{0}-M-(r_{+}-M)\cos\theta\right]}\;,$ (37)
where an integral constant $\frac{qr_{+}}{r_{0}}\frac{r_{0}-M}{r_{+}-M}$ has been selected to eliminate the divergence on the pole. The above equation (37) refers to the Coulomb field generated by a single static point charge $q$ in the Reissner-Nordström black hole spacetime. It is clear that the gauge potential vanishes at the extremal limit $r_{+}\rightarrow M$, which is an indication of the black hole Meissner effect Wald:1974np ; King:1975tt ; Bicak:1980du ; Bini:2008iwy .
@@
$\Delta A_{\theta}=-\frac{pr_{+}((r_{+}-M))\left(M-2r_{0}+(r_{+}-M)\cos\theta\right)\sin\theta}{r_{0}^{2}\sqrt{g_{rr}}|_{r=r_{0}}\left[r_{0}-M-(r_{+}-M)\cos\theta\right]^{2}}\;.$ (39)
@@
This result along with Eq.(28) gives the exact expression for the expected number of the entangling photon generated by the charged particle in the spatial separated quantum superposition state in the Reissner-Nordström black hole.
@@
6 Comparing the decoherence of Reissner-Nordström black holes with ordinary material systems
Biggs and Maldacena have found that the decoherence effect of a black hole is comparable to that of an ordinary material system Biggs:2024dgp . In this section, let us address the question of whether the decoherence caused by a Reissner-Nordström black hole is equivalent to that of an ordinary material system with the same temperature and charge.
To make a qualitative comparison, we consider the case where the charged particle is placed far from the horizon. In this scenario, the decohering flux presented in Eq.(40) can be approximated in the large $r_{0}$ limit as
$C=\frac{8p^{2}(M^{2}-Q^{2})r_{+}^{2}}{3\pi r_{0}^{6}}\;.$ (41)
Then the entangling photon number is approximated as
$\langle N\rangle=\frac{8p^{2}(M^{2}-Q^{2})^{3/2}}{3\pi r_{0}^{6}}\Delta T\;,$ (42)
and the constant decohering rate can be given by
$\Gamma=\frac{8p^{2}(M^{2}-Q^{2})^{3/2}}{3\pi r_{0}^{6}}\;.$ (43)
We first make a qualitative comparison with the local calculation of decohering effect presented in Biggs:2024dgp . It was shown that the decohering rate can be given by
$\Gamma\sim\frac{p^{2}}{r_{0}^{6}}S(\omega=0)\;,$ (44)
where $S(\omega)$ is the Wightman function. For the electromagnetic case, the Wightman function is related to the classical absorption cross section by the following relation Biggs:2024dgp
$S(\omega)\sim T_{H}\sigma_{\textrm{abs}}\;,$ (45)
where $T_{H}=\kappa_{+}/2\pi$ is the Hawking temperature of the outer horizon. The absorption cross section of the electromagnetic waves for the Reissner-Nordström black hole was computed by Gubser in Gubser:1997cm (see also Crispino:2009zza ), which is given by
$\sigma_{\textrm{abs}}=\frac{(r_{+}-r_{-})^{3}}{3T_{H}}\left(1+\frac{\omega^{2}}{4\pi^{2}T_{H}^{2}}\right)\;.$ (46)
It is straightforward to verify that the result from the local calculation is consistent with our result in Eq.(43).
According to the fluctuation dissipation theorem Kubo:1966fyg , the Wightman function is related to the response function $\chi(\omega)$ in the low frequency limit by the following relation
$\textrm{Im}\chi(\omega)\sim\frac{1}{2}\beta\omega S(\omega)\;,$ (47)
where $\beta$ is the inverse Hawking temperature. It is well known that the response function $\chi(\omega)$ describes the expectation value of an operator due to a small perturbation by an external source that couples to that operator. For the Reissner-Nordström black hole interacting with an external electric field, one can get the response function as
$\textrm{Im}\chi_{\textrm{BH}}(\omega)\sim\omega r_{+}^{2}(M^{2}-Q^{2})\sim\omega r_{+}^{4}\left(1-\frac{2Q^{2}}{r_{+}^{2}}\right)\;,$ (48)
where in the last step we have used the approximation condition $Q/r_{+}\ll 1$.
We now compare this with the response function of a solid conducting sphere with radius $R$, charge $Q$ and resistivity $\rho$. When the conducting sphere is placed in a static external electric field, the free charge $Q$ and the induce charge $Q_{i}$ will rearrange to cancel the electric field inside the sphere. The dipole moment due to the induced charge $Q_{i}$ can be approximated as $P\sim Q_{i}R$. However, we are mainly concerned with the case where the external electric field $\vec{E}_{\textrm{ext}}$ oscillates with the frequency $\omega$. In this scenario, the cancellation is incomplete due to the finite resistivity, which prevents charges from rearranging instantaneously Biggs:2024dgp . We can get
$\vec{E}_{\textrm{ext}}-\vec{E}_{i}=\rho\vec{J}\;,$ (49)
where $E_{i}$ is the electric field generated by the induced charge and $\vec{J}$ is the current density. Here, we ignore the electric field generated by the free charge $Q$.
The current density $\vec{J}$ can be decomposed into two parts, $\vec{J}_{Q}$ and $\vec{J}_{i}$, which correspond to the currents due to the movement of the free charge $Q$ and the induced charge $Q_{i}$ under the external electric field. The current density $\vec{J}_{i}$ can be related to the dipole moment via the relation $\int d^{3}x\vec{J}=\frac{d\vec{P}}{dt}$. This leads to the expression $J_{i}\sim\frac{-i\omega P}{R^{3}}$.
To estimate the current density $J_{Q}$, we recall the definition of the current density as
$J_{Q}=\frac{\Delta Q}{\Delta S\Delta t}\sim\frac{i\lambda\omega Q^{2}E_{\textrm{ext}}}{R^{2}}\;,$ (50)
where $\lambda$ is a numerical factor that may be determined by comparing with the black hole case. In this estimation, we have used the fact that the free charge $\Delta Q$ passing through the area element $\Delta S$ in the time interval $\Delta t$ should be proportional to the total free charge $Q$ and the electric field force $QE_{\textrm{ext}}$. Substituting these into Eq.(49), one can obtain
$E_{\textrm{ext}}\left(1-\frac{i\lambda\omega\rho Q^{2}}{R^{2}}\right)\sim\frac{P}{R^{3}}\left(1-i\omega\rho\right)\;.$ (51)
Then, we have
$\textrm{Im}\chi_{\textrm{cond}}(\omega)\sim\omega\rho R^{3}\left(1-\frac{\lambda Q^{2}}{R^{2}}\right)\;.$ (52)
We assume that the solid conducting sphere has the same size as the black hole, i.e. $R\sim r_{+}$. Compare Eq.(52) with the response function of the Reissner-Nordström black hole given in Eq.(48), we can see that $\rho/R$ should be of order unity. If $Q/R\ll 1$, the numerical factor $\lambda$ can be fixed to be $2$. Therefore, we can conclude that the decoherence effect caused by the Reissner-Nordström black hole is equivalent to that of an ordinary matter system with the same size and charge. This observation extends the conclusion of Biggs:2024dgp , which suggests that the decoherence due to a black hole horizon is equivalent to that of any other quantum system with the same temperature, to the charged case.
@@ 7 Conclusion and discussion
In summary, we have studied the DSW decoherence effect for a static charged particle in the Reissner-Nordström black holes. The explicit expression for the decoherence is obtained by evaluating the expected number of the entangling photon at late times. Our calculation is limited to the case of electromagnetic radiation from the charged particle in a spatial separated quantum superposition state. It is shown that the quantum superposition is decohered by the low frequency photons that propagate through the black hole horizon. We also discuss the case of the extremal Reissner-Nordström black hole. It is shown that the decoherence of quantum superposition is completely suppressed due to the black hole Meissner effect. We also compare the decoherence effect due to the Reissner-Nordström black hole with that of the ordinary material system and find that the decoherence effect caused by the Reissner-Nordström black hole is equivalent to that of an ordinary matter system with the same size and charge.
@@
Appendix A Calculation of “Coulomb” gauge potential on the horizon
The “Coulomb” gauge potential on the horizon, given by Eq.(37), is obtained by solving the Maxwell equations on the horizon, Eq.(35) and Eq.(36). Since the static charge is placed on the $z$ axis of a Reissner-Nordström black hole, the“Coulomb” gauge potential on the horizon is independent of the angular coordinate $\phi$. It is evident that the right-hand sides of Eq.(35) and Eq.(36) are gauge invariant. Thus, they can be calculated by using the gauge potential expressed in Eq.(32). However, this gauge potential have only the $t$ component. So the right-hand side of Eq.(36) is zero, which gives the fact that $A^{c}_{\phi}$ is a constant. Without loss of generality, we can set this constant to zero. The only non-vanishing component of the “Coulomb” gauge potential $A^{c}_{A}$ on the horizon is $A_{\theta}^{c}$.
For the right-hand side of Eq.(35), it can be shown that
$\frac{1}{2}\epsilon^{AB}\ \!{}^{*}\!F^{c}_{AB}=\frac{1}{\sqrt{q}}{}^{*}\!F_{\theta\phi}=\frac{\sqrt{-g}}{\sqrt{q}}F^{tr}=\partial_{r}A_{t}\;.$ (53)
By evaluating the derivative of $A_{t}$ on the horizon, one can get
$\frac{1}{2}\epsilon^{AB}\ \!{}^{*}\!F^{c}_{AB}=-\frac{q(r_{+}-M)\left[(r_{+}-M)\left(\cos^{2}\theta+1\right)-2(r_{0}-M)\cos\theta\right]}{r_{0}r_{+}\left[r_{0}-M-(r_{+}-M)\cos\theta\right]^{2}}\;.$ (54)
Finally integrating Eq.(35), one can get
$A_{\theta}^{c}=\frac{qr_{+}(r_{+}-M)\sin\theta}{r_{0}\left[r_{0}-M-(r_{+}-M)\cos\theta\right]}-\frac{qr_{+}(r_{0}-M)}{r_{0}(r_{+}-M)\sin\theta}+\frac{\lambda}{\sin\theta}\;,$ (55)
where $\lambda$ is a integral constant. It is obvious that the divergence of $A_{\theta}^{c}$ on the pole $\theta=0,\pi$ can be eliminated by selecting the integral constant $\lambda$ to be $\frac{qr_{+}(r_{0}-M)}{r_{0}(r_{+}-M)}$.